# InfoMatch: Entropy Neural Estimation for Semi-Supervised Image Classification

### v1

@@ 4 InfoMatch
Utilizing a neural network $\sigma(c_{j}|{\bm{x}}_{i},\theta)$ to approximate the information entropy for semi-supervised image classification, we have named our method InfoMatch. For each unlabeled data point ${\bm{x}}\in{\mathcal{X}}_{u}$, we conduct augmentation processing, creating one weak view, i.e., performing random flip, and two strong augmented views Cubuk et al. (2020) represented by ${\bm{x}}^{\prime}$, ${\bm{x}}^{(1)}$, and ${\bm{x}}^{(2)}$, respectively.
@@ 4.1 Minimize Upper Bound
According to Eq. (8), minimizing the cross entropy between the predicted probability distribution and the ground-truth distribution amounts to minimizing the likelihood function. By employing optimization algorithms like gradient descent, InfoMatch is gradually steered away from the incorrect distribution towards the accurate one.
@@
Additionally, we use CutMix to generate a new strong augmentation view on the top of weak augmentation, aiming to create more challenging samples and compel InfoMatch to fully extract meaningful features. To enhance computational efficiency, we shuffle the dataset in batches and randomly select an image, denoted as ${\bm{x}}^{\prime}_{r}$, at the corresponding position of ${\bm{x}}^{\prime}_{i}$. Subsequently, these images are used to generate CutMix images in a one-to-one manner. Furthermore, we evaluate the quantity of semantic information retained from the original images within the CutMix images, according to the size of the region. Consequently, the CutMix image and its corresponding label is expressed as follows:
@@ 4.2 Maximize Lower Bound
Inspired by BYOL Grill et al. (2020); Tian et al. (2023), we compute the contrastive loss by utilizing the positive pairs only, guaranteeing that samples of the same class are situated adjacent to each other in the embedding space. Furthermore, latent features often serve as covert structures or patterns derived from the data, offering a deeper understanding of the relationships within the data than merely predicting probability distributions. Consequently, we utilize the mutual information between the latent features corresponding to the two augmented views to approximate the lower bound of entropy.
@@
and the similarity defined in Eq. (16) implies that if any two logits are identical, the similarity is 1; otherwise, it tends to 0 as their distance increases.
@@
$\mathcal{L}_{\rm lower}={\color[rgb]{0,0,0}\frac{1}{n_{u}}}\sum_{i=1}^{n_{u}}\bigl{\|}{\bm{z}}^{(1)}_{i}-{\bm{z}}_{i}^{(2)}\bigr{\|}^{2}_{2}\,.$ (17)
@@ 4.3 InfoMatch Algorithm
In InfoMatch algorithm, we initiate with pseudo supervision and weak-to-strong strategies. This allows us to transform the entropy upper bound into pseudo-supervised cross entropy loss, serving as a supervisory signal for unlabeled data. Furthermore, to enrich the diversity of our dataset, we introduce a new strong augmentation method, CutMix, that enhances its generalization capabilities by formulating corresponding loss. Subsequently, we consider the potential feature contrastive loss between the two augmented views as a lower bound for dataset entropy. By maximizing this lower bound and minimizing the previously defined upper bound, we gradually optimize InfoMatch, even when limited labeled data is available.
$\mathcal{L}=\mathcal{L}_{\rm upper}+\lambda\mathcal{L}_{\rm lower}$ (18)
@@
1:  Initialization: $epoch=0$, $epoch_{max}$, and $\theta$.
2:  while $epoch\leq epoch_{max}$ do
@@
[…]
@@ 5.1 Experimental Setup
We evaluate InfoMatch on well-known benchmark datasets, including CIFAR-10/100 Krizhevsky and Hinton (2009), SVHN Netzer et al. (2011), and STL-10 Coates et al. (2011). Additionally, we conduct comprehensive experiments using varying amounts of labeled data.
To ensure a fair comparison, we follow the experimental setup used in Sohn et al. (2020); Zhang et al. (2021); Wang et al. (2023). Specifically, we employ standard stochastic gradient descent algorithm with cosine learning rate decay as the optimizer across all datasets, with an initial learning rate of 0.03 and a momentum of 0.9. For all experiments, we set the total number of iterations to $2^{20}$. InfoMatch performance is then evaluated using the EMA with a parameter of 0.999. Additionally, we utilize the Wide ResNet-28-2 Zagoruyko and Komodakis (2016) networks for CIFAR-10 and SVHN datasets, Wide ResNet-28-8 for CIFAR-100, and Wide ResNet-37-2 Zhou et al. (2020) for STL-10.
To evaluate the impact of the four distinct losses in our proposed approach, we introduce three additional non-negative hyperparameters. To ensure an equitable representation of both real and pseudo labels in the likelihood function and maintain equilibrium between RandAugment Cubuk et al. (2020) and CutMix Yun et al. (2019). Subsequently, we adjust the parameter $\lambda$ that regulates the entropy bounds to $0.002$.
In addition, we employ distinct threshold selection strategies to eliminate the impact of pseudo label errors on InfoMatch. For SVHN, we adopt a fixed threshold of 0.95. Conversely, for CIFAR-10/100 and STL-10 datasets, we utilize the self-adaptive thresholding method proposed by FreeMatch, which measures the overall learning progress by utilizing the expectation of the highest confidence across all unlabeled data in the current batch as the global threshold and assess the class-specific learning status by calculating the average prediction probabilities corresponding to each class.
Finally, to guarantee a precise and unbiased evaluation, we conducted multiple training sessions for each model using various random seeds and calculated the mean and standard deviation of the optimal accuracy achieved.
@@ 5.2 Main Results
We compare InfoMatch with full-supervised learning method and a range of representative semi-supervised learning methods, including pseudo-label-based methods such as FlexMatch Zhang et al. (2021) and FreeMatch Wang et al. (2023), as well as graph-based methods like SimMatchZheng et al. (2022) and SimMatchv2 Zheng et al. (2023). The top-1 error rates for CIFAR-10/100, SVHN, and STL-10 under various labeled data sizes are presented in Tab. 1 and Tab. 3.
Based on these results, it is evident that InfoMatch exhibits superior performance across all benchmarks, enhancing performance by an average of $1.49\%$, $0.55\%$, $0.13\%$, and $2.12\%$ on the four datasets. In particular, InfoMatch impressively lower the mean top-1 error rate from $8.07\%$ to $4.39\%$ ($-3.68\%$) in the CIFAR-10 dataset, which contains just 10 labeled data. While in in the STL-10 dataset with 40 labeled data, the top-1 error rate is reduces from $13.74\%$ to $9.86\%$ ( $-3.88\%$). Additionally, on both CIFAR-10 and SVHN datasets, our approach achieves remarkable performance improvements, surpassing not only other semi-supervised baselines but also outperforming fully-supervised learning methods across all benchmarks. Furthermore, compared to other methods, InfoMatch exhibits a smaller standard deviation across multiple experiments with varying seeds, thereby highlighting the superior stability and robustness of InfoMatch.
@@
(a) Top-1 accuracy
(b) Utilization of unlabeled data
Figure 1: Compare InfoMatch with FixMatch, FlexMatch and FreeMatch on the CIFAR-10 dataset with 40 labeled data in terms of top-1 accuracy and utilization of unlabeled data.
We assess InfoMatch on ImageNet to show its efficacy. Following the settings of FreeMatch, we select 100 labeled samples per class. Table 2 illustrates a comparison of the average top-1 and top-5 errors across different models. With identical parameters, InfoMatch hass a significant improvement of $4.49\%$ in top-1 accuracy and $2.86\%$ in top-5 accuracy over FreeMatch.
Table 2: Error rates (%) on ImageNet with 100 labels per class.
Top-1 Top-5
FixMatch Sohn et al. (2020) $43.66$ $21.80$
FlexMatch Zhang et al. (2021) $41.85$ $19.48$
FreeMatch Wang et al. (2023) $40.57$ $18.77$
InfoMatch (Ours) $\bm{36.21}$ $\bm{15.91}$
5.3 Ablation Study
InfoMatch primarily explores how to utilize of unlabeled data, where we use the weak-to-strong loss calculated based on two strong augmentation methods, RandAugment Cubuk et al. (2020) and CutMix Yun et al. (2019), represented by $\mathcal{L}^{u}_{\rm upper}$ and $\mathcal{L}^{c}_{\rm upper}$ respectively, as well as the contrastive loss $\mathcal{L}_{\rm lower}$ between two strong augmented views. To evaluate the impact of these losses, we conduct in-depth ablation studies on the CIFAR-10 dataset with 40 labels.
To determine the effectiveness of each loss in the framework, we conducted experiments by excluding some components. As shown in Fig. LABEL:fig:ablation, the results indicate that eliminating certain losses lead to different levels of decline in model performance. This suggests that each component is important and complementary.
Compared with RandAugment Cubuk et al. (2020), CutMix Yun et al. (2019) employs a stronger augmentation by combining two images, addressing the limitations of low mask-based augmentation ratio. This approach enables InfoMatch to more effectively comprehend and capture the essential features, thus enhancing the generalization performance of InfoMatch. However, utilizing CutMix exclusively may distort the original features, potentially leading InfoMatch to learn incorrect or distorted features during training, ultimately resulting in performance deterioration.
[…]
@@ 5.4 Qualitative Analysis
Initially, a noteworthy advantage of InfoMatch is its excellent convergence speed, which is illustrated in Fig. 1a. Within 100 epochs, InfoMatch has achieved a $94.70\%$ accuracy rate. By the time it reaches 200 epochs, our accuracy even exceeds the highest accuracy achieved by FreeMatch, reaching $95.49\%$. Additionally, as shown in Fig. 1b, in contrast to other methods, the utilization rate of unlabeled data in InfoMatch quickly rises to and stabilizes above $90\%$ after a brief decline, further encouraging InfoMatch to incorporate more unlabeled samples into training during the early stages.
@@ 6 Conclusion
We present an extension of the entropy neural estimation method to the semi-supervised domain, enabling to measure the uncertainty of classification models from an entropy perspective. We introduce an upper bound on the entropy of the ground-truth posterior, indicating that maximizing the likelihood function of InfoMatch corresponds to minimizing the upper bound on the posterior entropy. Similarly, we establish a lower bound on entropy for a given dataset, demonstrating that maximizing the mutual information between different augmented views of the data aligns with maximizing the lower bound of entropy. Drawing upon these two theories, we introduce InfoMatch to optimize it from two perspectives. To leverage unlabeled data more effectively, we employ strategies like weak-to-strong and pseudo supervision and introduce CutMix as a new strong augmentation view. Using optimization methods such as gradient descent, InfoMatch ensures that its predicted probability distribution gradually aligns with the ground-truth distribution, even in the presence of input perturbations. The results from various SSL benchmark tests demonstrate that InfoMatch achieves leading performance, especially in scenarios with limited labels, surpassing even fully-supervised learning methods. The experimental results strongly confirm the effectiveness of InfoMatch.

### v3

@@ 4 InfoMatch
Utilizing a neural network $\sigma(c_{j}|{\bm{x}}_{i},\theta)$ to approximate the information entropy for semi-supervised image classification, we call our method InfoMatch. For each unlabeled data point ${\bm{x}}\in{\mathcal{X}}_{u}$, we conduct augmentation processing, creating one weak view, i.e., performing random flip, and two strong augmented views Cubuk et al. (2020) represented by ${\bm{x}}^{\prime}$, ${\bm{x}}^{(1)}$, and ${\bm{x}}^{(2)}$, respectively.
@@ 4.1 Minimize Upper Bound
According to Eq. (8), minimizing the cross entropy between the predicted-probability distribution and the ground-truth distribution amounts to minimizing the likelihood function. By employing optimization algorithms like gradient descent, InfoMatch is gradually steered away from the incorrect distribution towards the accurate one.
@@
Additionally, we use CutMix to generate a new strong augmentation view on the top of weak augmentation, aiming to create more challenging samples and compel InfoMatch to fully extract meaningful features. To enhance computational efficiency, we shuffle the dataset in batches and randomly select an image, denoted as ${\bm{x}}^{\prime}_{r}$, at the corresponding position of ${\bm{x}}^{\prime}_{i}$. Subsequently, these images are used to generate CutMix images in a one-to-one manner. Furthermore, we evaluate the quantity of semantic information retained from the original images within the CutMix images, according to the size of the region. Consequently, the CutMix image and its corresponding pseudolabel is expressed by:
@@ 4.2 Maximize Lower Bound
Inspired by BYOL Grill et al. (2020), we compute the contrastive loss by utilizing the positive pairs only, guaranteeing that samples of the same class are situated adjacent to each other in the embedding space. Furthermore, latent features often serve as covert structures or patterns derived from the data, offering a deeper understanding of the relationships within the data than merely predicting probability distributions. Consequently, we utilize the mutual information between the latent features corresponding to the two augmented views to approximate the lower bound of entropy.
@@
and the similarity defined in Eq. (16) implies that if any two logits are identical, the similarity is one; otherwise, it tends to zero as their distance increases significantly.
@@
$\mathcal{L}_{\rm lower}=\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}\bigl{\|}{\bm{z}}^{(1)}_{i}-{\bm{z}}_{i}^{(2)}\bigr{\|}^{2}_{2}\,.$ (17)
@@ 4.3 InfoMatch Algorithm
In InfoMatch algorithm, we initiate with pseudo supervision and weak-to-strong strategies. This allows us to transform the entropy upper bound into pseudo-supervised cross entropy loss, serving as a supervisory signal for unlabeled data. Furthermore, to enrich the diversity of our dataset, we introduce a new strong augmentation method, CutMix, that enhances its generalization capabilities by formulating corresponding loss. Subsequently, we consider the potential feature contrastive loss between the two augmented views as a lower bound for dataset entropy. By maximizing this lower bound and minimizing the upper bound, we gradually optimize InfoMatch, even when limited labeled data is available.
@@
1:  Initialization: $epoch=0$, $epoch_{\max}$, and $\theta$.
2:  while $epoch\leq epoch_{\max}$ do
@@
$\mathcal{L}=\mathcal{L}_{\rm upper}+\lambda\mathcal{L}_{\rm lower}$ (18)
@@ 5.1 Experimental Setup
We evaluate InfoMatch on well-known benchmark datasets, including CIFAR-10/100 Krizhevsky and Hinton (2009), SVHN Netzer et al. (2011), STL-10 Coates et al. (2011), and ImageNet Deng et al. (2009). Additionally, we conduct experiments using varying amounts of labeled data.
To ensure a fair comparison, we follow the experimental setup as in FixMatch Sohn et al. (2020) and FreeMatch Wang et al. (2023). Specifically, we employ standard stochastic gradient descent algorithm with cosine learning rate decay as the optimizer across all datasets, with an initial learning rate of 0.03 and a momentum of 0.9. For all experiments, we set the total number of iterations to $2^{20}$. InfoMatch performance is then evaluated using the EMA with a parameter of 0.999. Additionally, for ImageNet, we maintain a batch size of 128 for both labeled and unlabeled samples, i.e., $n^{b}_{l}=n^{b}_{u}=128$, and utilize the ResNet-50 architecture He et al. (2016). While for other datasets, we adjust the batch sizes to $n^{b}_{l}=64$ and $n^{b}_{u}=448$, and employ the Wide ResNet variants, such as Wide ResNet-28-2 Zagoruyko and Komodakis (2016) and Wide ResNet-28-8 Zhou et al. (2020).
To evaluate the impact of the two distinct terms of the loss Eq. (18) in InfoMatch, we introduce a non-negative hyperparameter $\lambda$. Then, we ensure an equitable representation of both real and pseudolabels in the likelihood function and maintain equilibrium between RandAugment Cubuk et al. (2020) and CutMix Yun et al. (2019). Subsequently, we adjust the parameter $\lambda$ that regulates the entropy bounds to $0.002$.
In addition, we employ distinct threshold selection strategies to eliminate the impact of pseudolabel errors on InfoMatch. For SVHN, we adopt a fixed threshold of 0.95. Conversely, for other datasets, we utilize the self-adaptive thresholding method proposed by FreeMatch, which measures the overall learning progress by utilizing the expectation of the highest confidence across all unlabeled data in the current batch as the global threshold and assess the class-specific learning status by calculating the average prediction probabilities corresponding to each class.
Finally, to guarantee a precise and unbiased evaluation, we conduct multiple training sessions for each model using various random seeds and calculate the mean and standard deviation of the optimal accuracy achieved.
@@ 5.2 Main Results
We compare InfoMatch with full-supervised learning method and a range of representative semi-supervised learning methods, including pseudolabel-based methods such as FlexMatch Zhang et al. (2021) and FreeMatch Wang et al. (2023), as well as graph-based methods like SimMatchZheng et al. (2022) and SimMatchV2 Zheng et al. (2023). The Top-1 error rates for CIFAR-10/100, SVHN, and STL-10 under various labeled data sizes are presented in Tables 1 and 2.
Based on these results, it is evident that InfoMatch exhibits superior performance across all benchmarks, enhancing performance by an average of $1.49\%$, $0.55\%$, $0.13\%$, and $2.12\%$ on the four datasets. In particular, InfoMatch impressively lower the mean Top-1 error rate from $8.07\%$ to $4.39\%$ ($-3.68\%$) in the CIFAR-10 dataset, which contains just 10 labeled data. While in in the STL-10 dataset with 40 labeled data, the Top-1 error rate is reduces from $13.74\%$ to $9.86\%$ ( $-3.88\%$). Additionally, on both CIFAR-10 and SVHN datasets, our approach achieves remarkable performance improvements, surpassing not only other semi-supervised baselines but also outperforming fully-supervised learning methods across all benchmarks. Furthermore, comparing to other methods, InfoMatch exhibits a smaller standard deviation across multiple experiments with varying seeds, thereby highlighting the superior stability and robustness of InfoMatch.
@@
We assess InfoMatch on ImageNet to show its efficacy. Following the settings of FreeMatch, we select 100 labeled samples per class. Table 3 illustrates a comparison of the average Top-1 and Top-5 errors across different models. With identical parameters, InfoMatch has a significant improvement of $4.49\%$ in Top-1 accuracy and $2.86\%$ in Top-5 accuracy over FreeMatch.
@@
[…]
Method Top-1 Top-5
FixMatch Sohn et al. (2020) $43.66$ $21.80$
FlexMatch Zhang et al. (2021) $41.85$ $19.48$
FreeMatch Wang et al. (2023) $40.57$ $18.77$
InfoMatch $\bm{36.21}$ $\bm{15.91}$
Table 3: Error rates (%) on ImageNet with 100 labels per class.
5.3 Ablation Study
InfoMatch primarily explores how to utilize of unlabeled data, where we use the weak-to-strong supervision based on two strong augmentation methods, RandAugment Cubuk et al. (2020) and CutMix Yun et al. (2019), represented by $\mathcal{L}^{u}_{\rm upper}$ and $\mathcal{L}^{c}_{\rm upper}$ respectively, as well as the contrastive loss $\mathcal{L}_{\rm lower}$ between two strong augmented views. To evaluate the impact of these losses, we conduct in-depth ablation studies on the CIFAR-10 dataset with 40 labels.
To determine the effectiveness of each loss term, we conduct experiments by excluding some components. As shown in Fig. 1, Results indicate that eliminating certain losses lead to different levels of decline in model performance. This suggests that each component is important and complementary.
Comparing to RandAugment Cubuk et al. (2020), CutMix Yun et al. (2019) employs a stronger augmentation by combining two images, addressing the limitations of low mask-based augmentation ratio. This approach enables InfoMatch to more effectively comprehend and capture the essential features, thus enhancing the generalization performance of InfoMatch. However, utilizing CutMix exclusively may distort the original features, potentially leading InfoMatch to learning incorrect or distorted features during training, ultimately resulting in performance deterioration.
Figure 1: Ablation studies on the Cifar-10 dataset with 40 labels.
Figure 2: Compare InfoMatch with FixMatch, FlexMatch, and FreeMatch on the CIFAR-10 dataset with 40 labeled data in terms of Top-1 accuracy and utilization of unlabeled data. (a) Top-1 accuracy, and (b) Utilization of unlabeled data.
@@ 5.4 Qualitative Analysis
Initially, a noteworthy advantage of InfoMatch is its excellent convergence speed, which is illustrated in Fig. 2(a). Within 100 epochs, InfoMatch achieves a $94.70\%$ accuracy rate. By the time it reaches 200 epochs, our accuracy even exceeds the highest accuracy achieved by FreeMatch, reaching $95.49\%$. Additionally, as shown in Fig. 2(b), in contrast to other methods, the utilization rate of unlabeled data in InfoMatch quickly rises to and stabilizes above $90\%$ after a brief decline, further encouraging InfoMatch to incorporate more unlabeled samples into training during the early stages.
@@ 6 Conclusion
We present an extension of the entropy neural estimation method to the semi-supervised domain, enabling to measure the uncertainty of classification models from an entropy perspective. We introduce an upper bound on the entropy of the ground-truth posterior, indicating that maximizing the likelihood function of InfoMatch corresponds to minimizing the upper bound on the posterior entropy. Similarly, we establish a lower bound on entropy for a given dataset, demonstrating that maximizing the mutual information between different augmented views of the data aligns with maximizing the lower bound of entropy. Drawing upon these two theories, we introduce InfoMatch to optimize it from two perspectives. To leverage unlabeled data more effectively, we employ strategies like weak-to-strong and pseudo supervision, and introduce CutMix as a new strong augmentation view. Using optimization methods such as gradient descent, InfoMatch ensures that its predicted probability distribution gradually aligns with the ground-truth distribution, even in the presence of input perturbations. We introduce InfoMatch that seamlessly integrates pseudo supervision, consistency regularization, and mixing strategies to improve its generalization capabilities. The results from various SSL benchmark tests demonstrate that InfoMatch achieves good performance, especially in scenarios with limited labels, surpassing even fully-supervised learning methods. The experimental results confirm the effectiveness of InfoMatch.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant No. 62176108 , Natural Science Foundation of Qinghai Province of China under No. 2022-ZJ-929, Fundamental Research Funds for the Central Universities under No. lzujbky-2022-ct06, and Supercomputing Center of Lanzhou University. The corresponding author of the paper is Kun Zhan.